One card is selected at random from a deck of cards. Determine the probability that the card is a 3 or 7.
Answer by edjones(8007)

A standard deck has 52 cards.
There are 4 threes and 4 sevens.
4+4=8
8/52=2/13 the probability that the card is a 3 or 7.
.
Ed